get the value of x and y
x-y=5
3x+y=3

Answer by nerdybill(7384)

Applying the substitution method:
x-y=5 (equation 1)
3x+y=3 (equation 2)
.
Solve equation 1 for x:
x-y=5
x = y+5
Substitute the above into equation 2 and solve for y:
3x+y=3
3(y+5)+y=3
3y+15+y=3
4y+15=3
4y = -12
y = -3
.
Substitute the above into equation 1 and solve for x:
x-y=5
x-(-3)=5
x+3 = 5
x = 2
.
Solution:
x=2 and y=-3